What is the pH of an aqueous solution with a hydrogen ion concentration of [H+]=5.2×10−9 M?


1
Expert's answer
2022-05-02T17:02:07-0400

Given:

[H+] = 5.2×10−9 M


Solution:

pH = −log[H+] = −log(5.2×10−9) = 8.28

pH = 8.28 = 8.3 (2 sig figs)


Answer: pH = 8.3